# Dévissage for generation in derived categories

### v4

@@ 3. Results
The hypothesis tells us $E$ is a retract of an object $K$ which belongs to $\big{(}\mathbb{R}\pi_{\ast}D^{-}_{\operatorname{coh}}(Y)\big{)}^{\star n}$. Choose an integer $a$ such that $E$ belongs to $D^{>a}_{\operatorname{Qcoh}}(X)$. Consider the map $K\to L$ obtained from Lemma 3.9. We can check that $E$ is a retract of $L$ as the retraction $K\to E$ factors through $L$ for degree reasons. ∎
Theorem 3.11.
If $\pi\colon Y\to X$ is a proper surjective morphism of Noetherian schemes, then there is an integer $n$ such that $D^{b}_{\operatorname{coh}}(X)=\langle\mathbb{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle_{n}$.
@@
There is an integer $n$ such that each object of $D^{-}_{\operatorname{coh}}(X)$ belongs to $\langle\mathbb{R}\pi_{\ast}D^{-}_{\operatorname{coh}}(Y)\rangle_{n}$, see argument for [Aok21, Proposition 4.4]. The desired claim follows from Lemma 3.10. ∎
This gives the following immediate generalization of [Lan24, Theorem E] and [LO24, Lemma 3.15].
Corollary 3.12.
Let $\pi\colon Y\to X$ be a proper surjective morphism of Noetherian schemes. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(Y)$, then $\mathbb{R}\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$.
@@
Choose an integer $n$ such that $D^{b}_{\operatorname{coh}}(X)=\langle\mathbb{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle_{n}$, see Theorem 3.11. Let $E$ be an object of $D^{b}_{\operatorname{coh}}(X)$. We can find an object $E^{\prime}$ in $D^{b}_{\operatorname{coh}}(Y)$ such that $E$ is finitely built by $\mathbb{R}\pi_{\ast}E^{\prime}$. However, $\mathbb{R}\pi_{\ast}G$ finitely builds $\mathbb{R}\pi_{\ast}E^{\prime}$, and so, $E$ belongs to $\langle D^{b}_{\operatorname{coh}}(X)\rangle$. This completes the proof. ∎
@@
Let $X$ be a variety over a field. Recall that $D^{b}_{\operatorname{coh}}(X)$ has finite Rouquier dimension, see [Rou08, Theorem 7.38]. There exists a proper surjective morphism $\pi\colon\widetilde{X}\to X$ of varieties over $k$ such that $\dim\widetilde{X}=\dim X$ and $\widetilde{X}$ is a regular quasi-projective variety, cf. [dJ96, Theorem 4.1]. Let $\mathcal{L}$ be an ample line bundle on $\widetilde{X}$. By [Rou08, Proposition 7.31], we know that $G:=\bigoplus^{\dim X}_{i=-\dim X}\mathcal{L}^{\otimes i}$ is a strong generator for $D^{b}_{\operatorname{coh}}(\widetilde{X})$, and so $\mathbb{R}\pi_{\ast}G$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$.

### v5

@@ 3. Results
The hypothesis tells us $E$ is a retract of an object $K\in\big{(}\mathbf{R}\pi_{\ast}D^{-}_{\operatorname{coh}}(Y)\big{)}^{\star n}$. Choose an integer $a$ such that $E$ belongs to $D^{>a}_{\operatorname{Qcoh}}(X)$. Consider the map $K\to L$ obtained from Lemma 3.7. We can check that $E$ is a retract of $L$ as the retraction $K\to E$ factors through $L$ for degree reasons. ∎
Lemma 3.9.
If $\pi\colon Y\to X$ is a proper surjective morphism of Noetherian schemes, then there is an integer $n$ such that $D^{b}_{\operatorname{coh}}(X)=\langle\mathbf{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle_{n}$.
@@
There is an integer $n$ such that each object of $D^{-}_{\operatorname{coh}}(X)$ belongs to $\langle\mathbf{R}\pi_{\ast}D^{-}_{\operatorname{coh}}(Y)\rangle_{n}$. This can be observed from the argument for [Aok21, Proposition 4.4]. The desired claim follows from Lemma 3.8. ∎
We obtain generalizations of [Lan24, Theorem E], [LO25, Lemma 3.15], and [Gai13, Corollary 8.1.3].
Proposition 3.10.
Let $\pi\colon Y\to X$ be a proper surjective morphism of Noetherian schemes. If $G$ is a classical (resp. strong) generator for $D^{b}_{\operatorname{coh}}(Y)$, then $\mathbf{R}\pi_{\ast}G$ is a classical (resp. strong) generator for $D^{b}_{\operatorname{coh}}(X)$.
@@
We prove the case for classical generators as the other case follows similarly. By Lemma 3.9, there is an integer $n$ such that $D^{b}_{\operatorname{coh}}(X)=\langle\mathbf{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle_{n}$. Let $E$ be an object of $D^{b}_{\operatorname{coh}}(X)$. We can find an object $E^{\prime}$ in $D^{b}_{\operatorname{coh}}(Y)$ such that $E$ is finitely built by $\mathbf{R}\pi_{\ast}E^{\prime}$. However, $\mathbf{R}\pi_{\ast}G$ finitely builds $\mathbf{R}\pi_{\ast}E^{\prime}$, and so, $E$ belongs to $\langle D^{b}_{\operatorname{coh}}(X)\rangle$. This completes the proof. ∎
@@
[…]